Program A has a profit of $5,000 and an investment of $100,000, while program B has a profit of $10,000 and an investment of $220,000. Which program has the better ROI?
Answer #1

Project A

Return on Project A = Profit of Project A = $5000, Investment on Project A = $100,000

Return on Investment (ROI) of A = 5000/100000 = 0.05

Project B

Return on Project B = Profit of Project B = $10,000

Investment on Project B = $220,000

Return on Investment (ROI) of B = 10000/220000 = 0.04545455

ROI of Project A > ROI of Project B

Answer -> Project A has better ROI